# Beam Cross Sections Create Mixtures: Improving Feature Localization in Secondary Electron Imaging

### v1

@@ II-E SE Yield Estimation with Deterministic Number of Ions
$\mathrm{var}\!\left({\widehat{\eta}_{\mathrm{conv}}}\right)=\frac{1}{m}\left[(1-q)\eta_{1}+q\eta_{2}+q(1-q)(\eta_{2}-\eta_{1})^{2}\right].$ (15)
Furthermore, under a Poisson model as in (12), the sample mean is the maximum likelihood estimate (MLE) and the minimum-variance unbiased estimate of the distribution’s parameter.
Using the mixture model, we can improve the quality of the SE yield estimate. To derive a Cramér–Rao bound (CRB), we start by calculating the Fisher information (FI) about $q$ in $X$, with $\eta_{1}$ and $\eta_{2}$ assumed given, which we denote $\mathcal{I}_{X}(q;\eta_{1},\eta_{2})$.
From (9), the probability mass function (PMF) of $X$ is
By substitution, computing derivatives, and simplifying, one can find
$\mathcal{I}_{X}(q\,;\,\eta_{1},\eta_{2})$ $=\sum_{x=0}^{\infty}\frac{1}{x!}\frac{\left(\eta_{1}^{x}e^{-\eta_{1}}-\eta_{2}^{x}e^{-\eta_{2}}\right)^{2}}{(1-q)\eta_{1}^{x}e^{-\eta_{1}}+q\eta_{2}^{x}e^{-\eta_{2}}}.$ (17)
Since FI is additive, the reciprocal of $m\,\mathcal{I}_{X}(q\,;\,\eta_{1},\eta_{2})$ lower bounds the variance of an unbiased estimate for mixture parameter $q$ computed from the response to $m$ ions. Based on (10), one would form an estimate of SE yield $\eta$ from an estimate for $q$ through
$\widehat{\eta}=(1-\widehat{q})\eta_{1}+\widehat{q}\eta_{2}=\eta_{1}+(\eta_{2}-\eta_{1})\widehat{q}$. (18)
With the variance scaling implied by (18), we obtain
$\mathrm{var}\!\left({\widehat{\eta}}\right)\geq\frac{(\eta_{2}-\eta_{1})^{2}}{m\,\mathcal{I}_{X}(q\,;\,\eta_{1},\eta_{2})}.$ (19)
@@
Figure 3: Variances in SE yield estimation with and without exploiting a mixture model (9) as a function of mixing parameter $q$. Without the mixture model, SE yield is estimated with the sample mean; with the mixture model, an efficient estimate for $q$ is assumed to be rescaled according to (18). Plotted for $m=1$, $\eta_{1}=2$, $\eta_{2}=8$.
Figure 3 compares the variance (15) in SE yield estimation obtained without exploiting the mixture model with the lower bound in (19) obtained by using the mixture model. The sizable gap under the simplifying assumption of having a deterministic number of incident ions suggests that we will be able to improve SE imaging by using the mixture model (1). The reduction in variance comes in part from prior knowledge of $\eta_{1}$ and $\eta_{2}$; thus, the curves do not match even at $q=0$ or $q=1$, where the Poisson mixture reduces to a simple Poisson distribution.
@@ II-F Challenges
The calculations culminating in Fig. 3 isolate the effect of the beam cross section (randomness in transverse spatial position, see Fig. 1(c)) by ignoring both source shot noise and detector noise. The remainder of the paper studies settings where ion incidences are random in time (see Fig. 1(b)), with the aims of (i) finding theoretical limits of gains analogous to the gap in Fig. 3; and (ii) developing and demonstrating estimation methods. The processing of experimental data in Section V-D will include rudimentary consideration of detector noise.
@@
III Review of Time-Resolved Measurement
@@
In this section, we review three of the four measurement models from [25], omitting the discrete-time model that we do not build upon in this work; see [25, Sect. II] for additional details and illustrations. Then we summarize basic analyses for these models, slightly generalizing beyond the Poisson distribution for SE counts.
@@ III-A Measurement Models with Source Shot Noise
This modeling results in the following three measurement abstractions without detector noise:
@@
    This is an idealization of the standard operation of an SE imaging instrument.
  • Oracle: Observe
    $\left\{M,\,(T_{1},X_{1}),\,(T_{2},X_{2}),\,\ldots,\,(T_{M},X_{M})\right\}.$ (24)
    This data cannot be measured by any current instrument, but it serves a conceptual purpose.
@@
III-B Analyses of Conventional Observation
@@
$\lim_{\lambda\rightarrow\infty}\frac{\mathcal{I}_{Y}(\eta\,;\,\lambda)}{\lambda}=\frac{1}{\eta(1+\eta)}=\frac{1}{\eta}-\frac{1}{1+\eta},$ (30)
@@
III-C Estimators from Time-Resolved Measurements
From the oracle observation (24), one would naturally compute the sample mean of $\{X_{1},X_{2},\ldots,X_{M}\}$ to estimate $\eta$. The challenge in working with the TRM (25) is that $M$ is not available. Specifically, $M-\widetilde{M}$ is the number of $X_{i}$s that equal zero, and we do not observe these. For the case that $X_{i}$ is Poisson($\eta$) distributed, several estimators for $\eta$ are developed in [25, 26]. Here we summarize those with slight generalizations to non-Poisson distributions for $X_{i}$.
Since the sample mean $Y/M$ is an unbiased estimate with variance that vanishes with increasing dose, we may seek to use this but with some computable replacement for $M$. The thinning process described in Section III-A results in
$M\,|\,\widetilde{M}\sim\widetilde{M}+\mathop{\mathrm{Poisson}}(\lambda(1-\rho)),$ (31)
and thus $\mathrm{E}[\,{M\,|\,\widetilde{M}}\,]=\widetilde{M}+\lambda(1-\rho)$ [26]. Replacing $M$ with $\widetilde{M}+\lambda(1-\rho)$ results in
$\widehat{\eta}=\frac{Y}{\widetilde{M}+\lambda(1-\rho)},$ (32)
which is the MLE in the Poisson case. Note that $\rho$ is not known a priori; using (32) depends on a parametric form that relates $\rho$ to $\widehat{\eta}$, such as $\rho=1-e^{-\eta}$ when $X_{i}$ is Poisson($\eta$) distributed.
As introduced for the Poisson case in [26], since $M$ appears in the denominator of the sample mean, it is arguably more principled to multiply $Y$ by $\mathrm{E}[\,{1/M\,|\,\widetilde{M}}\,$] in an estimator. This is more complicated than (32) without having uniformly superior performance [26, Fig. 5]. In Section IV, we will thus concentrate on the MLE for the Poisson mixture derived in Section II-D.
III-D Analyses of Time-Resolved Measurement
As noted in Section III-B, FI computations require the distribution of $X_{i}$ to be explicit. We now review some results from [25, 26] for the Poisson case while highlighting extensions and distinctions for non-Poisson cases.
The estimators presented in Section III-C have $\{\widetilde{X}_{1},\widetilde{X}_{2},\ldots,\widetilde{X}_{\widetilde{M}}\}$ available but only use their sum $Y$. In the Poisson case, this simplification is related to the efficiency of the sample mean estimator and is explicitly justified in the derivation of the MLE. In Section IV, we will see that this does not carry over to Poisson mixtures.
@@ III-D Analyses of Time-Resolved Measurement
$\frac{1}{\lambda}\mathcal{I}_{\mathrm{TRM}}(\eta\,;\,\lambda)=\frac{1}{\lambda}\mathcal{I}_{\widetilde{M},\widetilde{T}_{1},\ldots,\widetilde{T}_{\widetilde{M}},\widetilde{X}_{1},\ldots,\widetilde{X}_{\widetilde{M}}}(\eta\,;\,\lambda)=\frac{1}{\eta}e^{-\eta}.$ (33)
The ratio between this and (30) is
@@
which varies from $1$ when $\eta=0$ to $\eta+1$ when $\eta$ is large. It represents the multiplicative information gain from TRM in the Poisson case. This gain is generalized to Poisson mixtures in Section IV.
An insightful decomposition from [26] can be partially generalized to non-Poisson distributions. Through steps detailed in [25],
@@
$\mathcal{I}_{\widetilde{M}}(\eta\,;\,\lambda)=\lambda\frac{e^{-\eta}}{e^{\eta}-1}$ (36)
@@
For the Poisson case of $\rho=1-e^{-\eta}$, (36) and (37) are equivalent through reparametrization [43, Ch. 2 (5.11)].
@@ IV Analyses for Poisson Mixture Distribution
In this section, we extend the results presented by Peng et al. [24, 25] to find the normalized Fisher information for the conventional observation (23) and time-resolved measurement (25) when the SE distribution $X$ is the two-component Poisson mixture (9). This combines the spatial and temporal aspects of the beam model in Section II-A, specialized for the sample in Section II-D. Our goals are to understand how the mixture-model gain suggested by Fig. 3 is modified by the inclusion of source shot noise and, similarly, how the TRM gain (34) is modified by the inclusion of beam cross-sectional mixing.
IV-A Analyses of Conventional Observation
@@
by substitution of (10) in (26) and (11) in (27). The validity of $\widehat{\eta}_{\rm baseline}$ as an unbiased estimator holds, and (29) specializes to
@@
Now to understand the theoretical limits for estimation of $Y$, we derive the FI about $q$ in $Y$. Instead of working purely algebraically with the PMF (16), let us derive the PMF of $Y$ by considering the mechanisms in the generation of $Y$. Given $M=m$ incident ions, they must be split as mim-i striking the sample where the SE yield is $\eta_{1}$ and $i$ striking where the SE yield is $\eta_{2}$ for some $i\in\{0,\,1\,\ldots,\,m\}$. Since the sum of independent Poisson random variables is a Poisson random variable, the result is a Poisson-distributed number of SEs with mean
@@
We now proceed to find the FI about $q$ in $Y$. Dependence on $q$ in $\mathrm{P}_{Y}(y)$ is only in $b_{m,i}$. Thus, we compute $b^{\prime}_{m,i}=\partialderivative{b_{m,i}}{q}$:
@@
$\mathcal{I}_{Y}(q\,;\,\eta_{1},\eta_{2},\lambda)=\sum_{y=0}^{\infty}\frac{(L(y))^{2}}{\mathrm{P}_{Y}(y)}.$ (47)
While (47) is not readily comprehensible, it can be used to numerically evaluate $\mathcal{I}_{Y}(q\,;\,\eta_{1},\eta_{2},\lambda)$.
Figure 4: Normalized Fisher information $\mathcal{I}_{Y}(q\,;\,\eta_{1},\eta_{2},\lambda)\,/\,\lambda$ as a function of dose $\lambda$ for $\eta_{1}=2$, $\eta_{2}=6$ and $q=0.6$. Also shown are the asymptotes (48) for low $\lambda$ and (49) for high $\lambda$.
At a coarse level, $\mathcal{I}_{Y}(q\,;\,\eta_{1},\eta_{2},\lambda)$ grows linearly with the dose $\lambda$, so it is plotted with normalization by $\lambda$ in Fig. 4. Importantly, the deviation from linearity is for normalized FI to be a decreasing function of $\lambda$. Since FI is additive, this proves that it is advantageous for a measurement at any dose to be replaced by a set of measurements with the same total dose. This is an argument in favor of time-resolved measurement, which approaches the behavior of infinitely many measurements, each with infinitesimal $\lambda$ [25].
@@
$\lim_{\lambda\rightarrow 0}\frac{\mathcal{I}_{Y}(q\,;\,\eta_{1},\eta_{2},\lambda)}{\lambda}=\sum_{y=1}^{\infty}\frac{1}{y!}\frac{\left(\eta_{1}^{y}e^{-\eta_{1}}-\eta_{2}^{y}e^{-\eta_{2}}\right)^{2}}{(1-q)\eta_{1}^{y}e^{-\eta_{1}}+q\eta_{2}^{y}e^{-\eta_{2}}}$ (48)
@@
$\lim_{\lambda\rightarrow\infty}\frac{\mathcal{I}_{Y}(q\,;\,\eta_{1},\eta_{2},\lambda)}{\lambda}=\frac{(\eta_{2}-\eta_{1})^{2}}{\eta+\eta^{2}+q(1-q)(\eta_{2}-\eta_{1})^{2}}$ (49)
are proven in the supplement. Remarkably, (48) differs from (17) only in the series starting at 1 rather than 0. So this normalized FI is less than $\mathcal{I}_{X}(q\,;\,\eta_{1},\eta_{2})$ by
@@
For comparisons to Fig. 3 and results in Section III, we would like the FI about $\eta$ rather than $q$. Scaling the FI according to the standard reparametrization formula [43, Ch. 2 (5.11)] here yields
$\mathcal{I}_{Y}(\eta\,;\,\eta_{1},\eta_{2},\lambda)=\frac{\mathcal{I}_{Y}(q\,;\,\eta_{1},\eta_{2},\lambda)}{(\eta_{2}-\eta_{1})^{2}}$ (51)
by using (10).
Applying the reparametrization to (49) gives
$\lim_{\lambda\rightarrow\infty}\frac{\mathcal{I}_{Y}(\eta\,;\,\lambda,\eta_{1},\eta_{2},\lambda)}{\lambda}=\frac{1}{\eta+\eta^{2}+q(1-q)(\eta_{2}-\eta_{1})^{2}}.$ (52)
The reciprocal of this FI matches the MSE (40) of the baseline estimator $\widehat{\eta}_{\rm baseline}$, which shows again—as in Section III-B—that $\widehat{\eta}_{\rm baseline}$ is efficient in the high $\lambda$ limit. Improved estimation at useful doses depends on time-resolved measurement.

### v2

@@ II-E Potential Improvement from Mixture Model
$\mathrm{var}\!\left({\widehat{\eta}_{\mathrm{SM}}}\right)=\frac{1}{m}\left[(1-q)\eta_{1}+q\eta_{2}+q(1-q)(\eta_{2}-\eta_{1})^{2}\right].$ (13)
The Cramér–Rao bound for estimation of $\eta$ without knowledge of $\eta_{1}$ and $\eta_{2}$, implicitly exploiting the mixture model if possible, is derived in Supplementary Note 1. It is numerically indistinguishable from (13).
When the values of $\eta_{1}$, and $\eta_{2}$ are known, they provide the constraint of $\eta\in[\eta_{1},\,\eta_{2}]$. In the absence of a model for the $X_{i}$ distribution, it is natural to base an estimate on $\widehat{\eta}_{\mathrm{SM}}$, possibly with some further computation. [^1] If $X_{i}$ is Poisson, $\widehat{\eta}_{\mathrm{SM}}$ is a sufficient statistic for estimation of $\eta$ from $(X_{1},\,X_{2},\,\ldots,X_{m})$. Estimation from the sum $\sum_{i=1}^{m}X_{i}$ is also justified under the conventional observation model of Section III-A. For $m\gtrsim$ 30, the central limit theorem implies that $\widehat{\eta}_{\mathrm{SM}}$ is approximately Gaussian, and thus clipping $\widehat{\eta}_{\mathrm{SM}}$ to the interval $\eta_{1},\,\eta_{2}$] gives an approximate constrained maximum likelihood estimate of $\eta$ from $X_{1}+X_{2}+\cdots+X_{m}$:
$\widehat{\eta}_{\mathrm{clip}}=\operatorname*{median}(\{\eta_{1},\,\widehat{\eta}_{\mathrm{SM}},\,\eta_{2}\}).$ (14)
Reduction in MSE due to clipping occurs for $q$ near 0 or near 1, as illustrated in Fig. 3 for $\eta_{1}=2$, $\eta_{2}=8$, and $m=100$.
@@
Figure 3: Variances in SE yield estimation with and without exploiting a mixture model (7) and knowledge of $(\eta_{1},\eta_{2})$ as a function of mixing parameter $q$. Without a distributional model, SE yield is estimated with the sample mean, which can be clipped to exploit knowledge of $(\eta_{1},\eta_{2})$. With the mixture model, an estimate achieving the Cramér–Rao bound would provide an improvement. Plotted for $m=100$, $\eta_{1}=2$, $\eta_{2}=8$.
In the case of $\eta_{1}$ and $\eta_{2}$ being known, we can use the mixture model to improve the quality of the SE yield estimate. To derive a Cramér–Rao bound (CRB), we start by calculating the Fisher information (FI) about $q$ in $X$, with $\eta_{1}$ and $\eta_{2}$ given, which we denote $\mathcal{I}^{\mathrm{mix}}_{X}(q;\eta_{1},\eta_{2})$. [^2] In our notation, $\mathcal{I}_{Z}^{\mathrm{model}}(\theta\,;\,\phi_{1},\phi_{2})$ is the Fisher information in observation $Z$ about parameter $\theta$ when parameters $\phi_{1}$ and $\phi_{2}$ are known. Since the SE yield parameter $\eta$ is applicable to multiple data models, the data model is emphasized with superscript text.
Using (8),
$\frac{\partial\log\mathrm{P}_{X}(x\,;\,q,\eta_{1},\eta_{2})}{\partial q}=\frac{-(\eta_{1}^{x}e^{-\eta_{1}}-\eta_{2}^{x}e^{-\eta_{2}})}{(1-q)\eta_{1}^{x}e^{-\eta_{1}}+q\eta_{2}^{x}e^{-\eta_{2}}}.$ (15)
Thus, by substitution and simplification,
$\mathcal{I}^{\mathrm{mix}}_{X}(q\,;\,\eta_{1},\eta_{2})$ $=-\mathrm{E}\!\left[\,{\left(\frac{\partial\log\mathrm{P}_{X}(X\,;\,q,\eta_{1},\eta_{2})}{\partial q}\right)^{2}}\,\right]$
$=\sum_{x=0}^{\infty}\frac{1}{x!}\frac{\left(\eta_{1}^{x}e^{-\eta_{1}}-\eta_{2}^{x}e^{-\eta_{2}}\right)^{2}}{(1-q)\eta_{1}^{x}e^{-\eta_{1}}+q\eta_{2}^{x}e^{-\eta_{2}}}.$ (16)
The FI about $\eta$ can be obtained with a well-known property of reparameterization[41, Ch. 2 (5.11)]. We divide by $(\partial\eta/\partial q)^{2}=(\eta_{2}-\eta_{1})^{2}$ to get
$\mathcal{I}^{\mathrm{mix}}_{X}(\eta\,;\,\eta_{1},\eta_{2})$ $=\frac{\mathcal{I}^{\mathrm{mix}}_{X}(q\,;\,\eta_{1},\eta_{2})}{(\eta_{2}-\eta_{1})^{2}}.$ (17)
Since FI is additive, the reciprocal of $m\,\mathcal{I}^{\mathrm{mix}}_{X}(\eta\,;\,\eta_{1},\eta_{2})$ lower bounds the variance of an unbiased estimate for $\eta$ computed from the response to $m$ ions. This CRB is included in Fig. 3. The sizeable gaps from the MSEs of $\widehat{\eta}_{\mathrm{SM}}$ and $\widehat{\eta}_{\mathrm{clip}}$ suggest that using the mixture model—with knowledge of $\eta_{1}$ and $\eta_{2}$—provides a potential advantage that is not merely due to the knowledge of $\eta_{1}$ and $\eta_{2}$. We seek to realize this advantage.
@@ II-F Challenges
The calculations culminating in Fig. 3 isolate the effect of the beam cross section (randomness in transverse spatial position, see Fig. 1(c)) from the effects of source shot noise and detector noise. The remainder of the paper studies settings where ion incidences are random in time (see Fig. 1(b)), with the aims of (i) finding theoretical limits of gains analogous to the gap in Fig. 3; and (ii) developing and demonstrating estimation methods. The processing of experimental data in Section V-D will include rudimentary consideration of detector noise.
@@
III Measurement Models
@@
In this section, we review two of the four measurement models from [25], omitting the oracle and discrete-time models that we do not build upon in this work; see [25, Sect. II] for additional details and illustrations. Then we summarize basic analyses for these models, slightly generalizing beyond the Poisson distribution for SE counts.
@@ III-A Measurement Models with Source Shot Noise
This modeling results in the following measurement abstractions without detector noise:
@@
    This is an idealization of the standard operation of existing SE imaging instruments.
@@
III-B Analyses of Conventional Observations
@@
$\lim_{\lambda\rightarrow\infty}\frac{\mathcal{I}^{\mathrm{Poisson}}_{Y}(\eta\,;\,\lambda)}{\lambda}=\frac{1}{\eta(1+\eta)}=\frac{1}{\eta}-\frac{1}{1+\eta},$ (27)
@@
III-C Analyses of Time-Resolved Measurements
As noted in Section III-B, FI computations require the distribution of $X_{i}$ to be explicit. We now review some results from [25, 26] for the Poisson case while highlighting a useful intermediate result that applies without regard to the $X_{i}$ distribution.
@@ III-C Analyses of Time-Resolved Measurements
$\frac{1}{\lambda}\mathcal{I}^{\mathrm{Poisson}}_{\mathrm{TRM}}(\eta\,;\,\lambda)=\frac{1}{\lambda}\mathcal{I}^{\mathrm{Poisson}}_{\widetilde{M},\widetilde{T}_{1},\ldots,\widetilde{T}_{\widetilde{M}},\widetilde{X}_{1},\ldots,\widetilde{X}_{\widetilde{M}}}(\eta\,;\,\lambda)=\frac{1}{\eta}-e^{-\eta}.$ (28)
The ratio between (28) and (27) is
@@
which varies from $1$ when $\eta=0$ to $\eta+1$ when $\eta$ is large. It represents the multiplicative information gain from TRM in the Poisson case. The analogous gain for Poisson mixtures is derived in Section IV.
An insightful decomposition from [26] is presented here because we will use it in Section IV. Through steps detailed in [25],
@@
$\mathcal{I}^{\mathrm{Poisson}}_{\widetilde{M}}(\eta\,;\,\lambda)=\lambda\frac{e^{-\eta}}{e^{\eta}-1}$ (31)
@@
For the Poisson case of $\rho=1-e^{-\eta}$, (31) and (32) are equivalent through reparametrization [41, Ch. 2 (5.11)] because $(\partial\eta/\partial\rho)^{2}=e^{2\eta}$.
@@ IV Analyses for Poisson Mixture Distribution
We now find the normalized Fisher information for the conventional observation (21) and time-resolved measurement (22) when the SE distribution $X$ is the two-component Poisson mixture (7). This combines the spatial and temporal aspects of the beam model in Section II-A, specialized for the sample in Section II-D. Our goals are to understand how the mixture-model gain suggested by Fig. 3 is modified by the inclusion of source shot noise and, similarly, how the TRM gain (29) is modified by the inclusion of beam cross-sectional mixing.
IV-A Analyses of Conventional Observations
@@
by substitution of (9) in (23) and (10) in (24); recall that $\eta=(1-q)\eta_{1}+q\eta_{2}$, making the second expression more compact. The validity of $\widehat{\eta}_{\rm baseline}$ as an unbiased estimator holds, and (26) specializes to
@@
Now to understand the theoretical limits for estimation of $q$ from $Y$, we derive the FI about $q$ in $Y$. Instead of working purely algebraically with the PMF (8), let us derive the PMF of $Y$ by considering the mechanisms in the generation of $Y$. Given $M=m$ incident ions, they must be split as mim-i striking the sample where the SE yield is $\eta_{1}$ and $i$ striking where the SE yield is $\eta_{2}$ for some $i\in\{0,\,1\,\ldots,\,m\}$. Since the sum of independent Poisson random variables is a Poisson random variable, the result is a Poisson-distributed number of SEs with mean
@@
We now proceed to find the FI about $q$ in $Y$. Dependence on $q$ in $\mathrm{P}_{Y}(y)$ is only in $b_{m,i}$. Thus, we compute $b^{\prime}_{m,i}={{\partial b_{m,i}}/{\partial q}}$:
@@
$\mathcal{I}^{\mathrm{mix}}_{Y}(q\,;\,\eta_{1},\eta_{2},\lambda)=\sum_{y=0}^{\infty}\frac{(L(y))^{2}}{\mathrm{P}_{Y}(y)}.$ (42)
While (42) is not readily comprehensible, it can be used to numerically evaluate $\mathcal{I}^{\mathrm{mix}}_{Y}(q\,;\,\eta_{1},\eta_{2},\lambda)$.
Figure 4: Normalized Fisher information $\mathcal{I}^{\mathrm{mix}}_{Y}(q\,;\,\eta_{1},\eta_{2},\lambda)\,/\,\lambda$ as a function of dose $\lambda$ for $\eta_{1}=2$, $\eta_{2}=6$ and $q=0.6$. Also shown are the asymptotes (43) for low $\lambda$ and (44) for high $\lambda$. The ratio of the asymptotes is the information gain factor $\beta_{\mathrm{mixture}}$ given in (59). It is the factor by which dose or MSE could be reduced due to the use of time-resolved measurement.
At a coarse level, $\mathcal{I}^{\mathrm{mix}}_{Y}(q\,;\,\eta_{1},\eta_{2},\lambda)$ grows linearly with the dose $\lambda$, so it is plotted with normalization by $\lambda$ in Fig. 4. Importantly, the deviation from linearity is for normalized FI to be a decreasing function of $\lambda$. Since FI is additive, this proves that it is advantageous for a measurement at any dose to be replaced by a set of measurements with the same total dose. This is an argument in favor of time-resolved measurement, which approaches the behavior of infinitely many measurements, each with infinitesimal $\lambda$ [25].
@@
$\lim_{\lambda\rightarrow 0}$ $\frac{\mathcal{I}^{\mathrm{mix}}_{Y}(q\,;\,\eta_{1},\eta_{2},\lambda)}{\lambda}=\sum_{y=1}^{\infty}\frac{1}{y!}\frac{\left(\eta_{1}^{y}e^{-\eta_{1}}-\eta_{2}^{y}e^{-\eta_{2}}\right)^{2}}{(1-q)\eta_{1}^{y}e^{-\eta_{1}}+q\eta_{2}^{y}e^{-\eta_{2}}}$ (43)
@@
$\lim_{\lambda\rightarrow\infty}$ $\frac{\mathcal{I}^{\mathrm{mix}}_{Y}(q\,;\,\eta_{1},\eta_{2},\lambda)}{\lambda}=\frac{(\eta_{2}-\eta_{1})^{2}}{\eta+\eta^{2}+q(1-q)(\eta_{2}-\eta_{1})^{2}}$ (44)
are proven in Supplementary Note 2. Remarkably, (43) differs from (16) only in the series starting at 1 rather than 0. So this normalized FI is less than $\mathcal{I}^{\mathrm{mix}}_{X}(q\,;\,\eta_{1},\eta_{2})$ by
@@
For comparisons to Fig. 3 and results in Section III, we would like the FI about $\eta$ rather than $q$. Analogously to (17), we divide by $(\partial\eta/\partial q)^{2}=(\eta_{2}-\eta_{1})^{2}$:
$\mathcal{I}^{\mathrm{mix}}_{Y}(\eta\,;\,\eta_{1},\eta_{2},\lambda)$ $=\frac{\mathcal{I}^{\mathrm{mix}}_{Y}(q\,;\,\eta_{1},\eta_{2},\lambda)}{(\eta_{2}-\eta_{1})^{2}}.$ (46)
Applying the same reparametrization to (44) gives
$\lim_{\lambda\rightarrow\infty}$ $\frac{\mathcal{I}^{\mathrm{mix}}_{Y}(\eta\,;\,\eta_{1},\eta_{2},\lambda)}{\lambda}=\frac{1}{\eta+\eta^{2}+q(1-q)(\eta_{2}-\eta_{1})^{2}}.$
The reciprocal of this normalized FI matches the MSE (35) of the baseline estimator $\widehat{\eta}_{\rm baseline}$, which shows again—as in Section III-B—that $\widehat{\eta}_{\rm baseline}$ is efficient in the high $\lambda$ limit. Improved estimation at useful doses depends on time-resolved measurement.